OpenStudy (anonymous):

I need help simplifying: a^(1/2)[a^(3/2)-2a^(1/2)]

OpenStudy (a_clan):

\[\sqrt{a}[a ^{3/2} - 2\sqrt{a}]\] ?

OpenStudy (a_clan):

\[=a ^{2}-2a\] \[=a(a-2)\]